Imagine you stopped working for uber in December 31st of 2017, which you used to make an average of $26,400 a year, to open your own coffee shop on January 1st. 2018. Your coffee shop employs two workers each earning $15 an hour, each working 140 hours a month per year. Your total annual cost of raw materials in 2018 was $4,800. In 2018 you Pais "Joe's remodeling Co." to remodel your store at a total cost of $5,300 Your coffee shop's monthly lease in 2018 was $1,200. In 2018 your shop made total revenue equal to $189,400. What is your explicit cost? what is your implicit cost? what is your total economic cost? what is your "accounting profit? what is your " economic profit?
Answer #1

Explicit cost = amount paid to employees + amount incurred on raw materials + amount incurred on remodeling + lease amount

= 2*15*140*12 + 4800 + 5300 + 1200*12

= 74900

Implicit cost = lost earnings = 26400

Revenue = 189400

Total economic cost = explicit cost + implicit cost = 74900 + 26400 = 101300

Accounting profit = revenue – explicit cost = 189400 – 74900 = 114500

Economic profit = revenue – total economic cost = 88100